# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to sequence sub-goals and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, many studies explore the use of foundation models to solve challenges like long-horizon problems and effectively exploring and designing reward functions.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. ROBOFUME (Yang et al. (2024a)) employs a fine-tuned VLM as the sparse reward function for the RL algorithm, tackling the issue of the extensive human supervision needed for training or fine-tuning a policy in the real world. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling action efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
@@ 8 Manipulation Data Generation
Demonstration data plays a crucial role in robotic manipulation, particularly in the context of imitation learning (Padalkar et al. (2023b)). A common approach for gathering such demonstrations is human teleoperation in the real world. However, collecting real-world data often requires significant human labor and specialized teleoperation equipment. Recently, there has been a growing number of excellent developments in low-cost teleoperation hardware, which enables the collection of high-quality demonstration data (Fang et al. (2024); Cheng et al. (2024)).
@@
To collect data in real environments, human effort is required for scene setup and data annotation (Sermanet et al. (2023)). There are currently two methods for data collection: the bottom-up approach and the top-down step-by-step approach. The bottom-up approach focuses on selecting a task to perform based on the current scene. Then, it uses methods like crowd-sourcing to label the data. The top-down approach involves a step-by-step process where decision-makers assign task labels and manage overheads, such as resets and scene preparations (Sermanet et al. (2023)). The robot then performs tasks according to these labels. RoboVQA (Sermanet et al. (2023)) shows that the bottom-up approach is more efficient in data collection compared to the top-down step-by-step approach. DIAL (Xiao et al. (2022a)) uses a fine-tuned CLIP to replace humans in labeling robot trajectories during bottom-up data collection. This transforms the robot manipulation dataset on the internet into the robot-language manipulation dataset. PAFF (Ge et al. (2023)) points out that incorrect robot trajectories can be linked to new tasks and uses fine-tuned CLIP to label the incorrect robot trajectories with appropriate task labels. The above methods demonstrate that high-level cognitive models can assist in data annotation. SOAR (Zhou et al. (2024b)) shows that integrating a high-level cognitive model with a low-level control policy can result in a fully autonomous data-collection system in varied real-world environments.
@@
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023c)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge.
@@
Compared to single-frame images and language data on the internet, internet videos contain information on the physics and dynamics of the world, as well as on human behaviors and actions (Chandrasegaran et al. (2024)). This information is precisely what is required for manipulation tasks. Therefore, in this section, we also introduce the internet-scale video data for robot learning.
@@
Aloha(Zhao et al. (2023a)) Online Teleoperation Joint Copy Third-person View \$2000(include robot) Dual-Arm; 2F Gripper
@@
GELLO(Wu et al. (2023c)) Online Teleoperation Joint copy Third-person View \$300 Dual-Arm; 2F Gripper
@@
Human Plus(Fu et al. (2024)) Online Teleoperation End-effectors mapping Third-person View \$30(A RGB Camera) Humanoid Robot; Dexterous Hand
@@
Transteleop(Li et al. (2020)) Online Teleoperation End-effectors mapping Third-person View \$170 Dual-Arm; Dexteours Hand
@@
OPEN TEACH(Iyer et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$500 Dual-Arm; Dexteours Hand/2F Gripper;
@@
Open-TeleVision(Cheng et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$3499 Humanoid Robot; Dexteours Hand
@@
AirExo (Fang et al. (2024)) Offline Teleoperation Joint copy - \$600 Dual-Arm; 2F Gripper
@@
UMI(Chi et al. (2024)) Offline Teleoperation End-effectors mapping - \$371 Dual-Arm; 2F Gripper
@@
DexCap(Wang et al. (2024c)) Offline Teleoperation End-effectors mapping - \$4000 Dual-Arm; Dexterous Hand
@@
VideoDex(Shaw et al. (2023)) Offline Teleoperation End-effectors mapping - - Single-Arm; Dexterous Hand
@@ 8.1 Low-cost Teleoperation Device
The current low-cost teleoperation can be categorized into two types: online teleoperation and offline teleoperation. The distinction is similar to the difference between SLAM and SFM. Online teleoperation is a closed-loop interaction between a demonstrator and a robot (Darvish et al. (2023)). In the forward process, human motion is measured using devices that combine various sensors, such as vision, IMUs, or multi-joint encoders. The motion data from the demonstrator is then retargeted to the robot’s space. This allows the robot to accurately follow the demonstrator’s demonstrated trajectory. During the backward process, sensor data from the robot, such as forces, torques, and tactile information, should be retargeted to the demonstrator’s space. As a result, the demonstrator can experience an immersive teleoperation environment by sensor data feedback. At the same time, the synchronization and real-time performance between the forward and backward processes are also crucial (Darvish et al. (2023)). Offline teleoperation remove the reliance on real robots during data collection compared to online teleoperation (Chi et al. (2024)). Demonstrators directly perform tasks using handheld or wearable devices (Fang et al. (2024); Chi et al. (2024); Wang et al. (2024c)) or using cameras to record the task execution process (Shaw et al. (2023)). They do not need to supervise real robots to complete the tasks and operate tasks using human’s direct view perspective. Therefore, offline teleoperation lacks the backward feedback process. Without relying on real robots, the devices become more portable and intuitive. However, this increases the precision requirements for the retargeting algorithm.
@@
The differences among current low-cost teleoperation devices lie primarily in two aspects. One is human motion measurement on both online teleoperation and offline teleoperation. The other is visual feedback on online teleoperation. Human motion measurement component can be broadly categorized into two classes: one aimed at capturing and mapping the pose of end-effectors (Cheng et al. (2024); Liu et al. (2022a); Fu et al. (2024); Chi et al. (2024); Li et al. (2020)), and one exploited devices for joint copy (Zhao et al. (2023a); Wu et al. (2023c); Fang et al. (2024)). Visual feedback can be generally classified into two types third-person view and first-person view (Cheng et al. (2024)). The third-person view shows the demonstrator from an external position, offering a broader perspective of surroundings. In contrast, the first-person view mimics the robot’s perspective, providing an immersive and realistic experience such as teleoperation with VR/AR headset.
@@
For approaches capturing and mapping the pose of end-effectors, the common low-cost capturing devices include SpaceMouse (Liu et al. (2022a); Zhu et al. (2023b)), cameras (Cheng et al. (2024); Fu et al. (2024); Iyer et al. (2024); Shaw et al. (2023); Li et al. (2019); Fang et al. (2020a)), VR controllers (De Pace et al. (2021); Nakanishi et al. (2020)) and IMU sensors (Chi et al. (2024); Li et al. (2020); Fang et al. (2017a, b)). The SpaceMouse based method passes the position and orientation of the SpaceMouse as action commands of end-effectors. This method is low-cost, easy operation, and easy implementation, but it is not suitable for dual-arm operations. In contrast, methods based on cameras and VR are well suited for bimanual teleoperation and VR offers the advantage of visual feedback compared to cameras. However, teleoperation methods based on cameras and VR heavily relies on the accuracy of pose estimation algorithms and often affected by occlusion (Fu et al. (2024); Pavlakos et al. (2024); Iyer et al. (2024); Cheng et al. (2024)). The main advantage of teleoperation devices based on IMU sensors lies in their wearability (Li et al. (2020); Chi et al. (2024); Wang et al. (2024c)). Due to this advantage, UMI (Chi et al. (2024)) and DexCap (Wang et al. (2024c)) develope wearable devices capable of in-the-wild teleoperation and offline data collection.
@@
Above systems work in cartesian space, which needs inverse kinematic (IK) solver and off-the-shelf IK often suffering from fails when operating near singularities of the robot. Although some bilateral teleoperation systems use haptic feedback to provide a tangible sense of the robot’s kinematic constraints, they do not address the challenges of very tight operational spaces (Silva et al. (2009)). Therefore, multi-joint encoder teleoperation devices can solve the IK problem by working in the joint space. The current design of multi-joint teleoperation devices is mainly divided into isomorphic and non-isomorphic devices (Wu et al. (2023c)). Isomorphic devices refer to teleoperation systems using standard servo-based robotic arms to control manipulators with similar size and kinematics (Zhao et al. (2023a)), while non-isomorphic devices use such arms to control manipulators with different size and kinematic properties. Non-isomorphic devices use kinematically equivalent structures based on DH parameters to map joint spaces between different properties (Wu et al. (2023c)). AirExo (Fang et al. (2024)) expands this low-cost and scalable platform into a wearable device to collect cheap in-the-wild demonstrations at scale.
@@
As for teleoperation visual feedback, most of methods (Liu et al. (2022a); Zhu et al. (2023b); Fu et al. (2024); Li et al. (2020); Zhao et al. (2023a); Wu et al. (2023c)) are use third side view that observe the robot task with the operator’s own eyes directly. However, this observation involves some visual errors. For example, there may be inaccuracies in the distance between the gripper and the object being manipulated. While for first-person view, due to wearing VR head (Cheng et al. (2024); Iyer et al. (2024); De Pace et al. (2021); Nakanishi et al. (2020)), it allows operators to perceive the robot’s surroundings immersively. However, long time to use VR headset can lead to fatigue.
@@ 8.2 Simulator
The current mainstream simulators (Zhou et al. (2023)) include PyBullet (Coumans and Bai (2016)), MuJoCo (Todorov et al. (2012)), CoppeliaSim (Rohmer et al. (2013)), NVIDIA Omniverse and Unity. Pybullet is easy to use and integrate, but its graphics are quite basic. It is not suitable for applications that require complex visual effects. Therefore, Pybullet is often used together with Blender (Shi et al. (2024)). Mujoco offers a high-precision physics engine. It is suitable for simulating articulated and deformable object manipulation. However, it has a high entry barrier for beginners. CoppeliaSim offers a wide range of ready-made environments, objects, and prototyping robotic systems for users. However, when dealing with many robots or complex scenes, CoppeliaSim may encounter performance issues. NVIDIA Omniverse provides real-time physics simulation and realistic rendering. However, it requires significant computational resources. NVIDIA Omniverse offers many interfaces. Users can use these to develop various applications. For example, Issac Gym is a platform for robot reinforcement learning, developed using Omniverse. Unity offers rich visual effects and a user-friendly interface. It allows for the creation of highly interactive applications. However, its physics engine is still not precise enough. The basic components of a simulator are the physics engine and the renderer. Improvements in these components can enhance the capability of sensors in simulations, such as optical tactile sensors (Chen et al. (2023d)). Learning-based simulators also show great potential. For example, Sora (Brooks et al. (2024)) and UniSim (Yang et al. (2023b)) use vast amounts of data from the internet to simulate the visual effects of many different actions.

### v7

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (kober2013reinforcement). Foundation models have demonstrated the emergence of common sense reasoning, the ability to sequence sub-goals and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (jin2024robotgpt) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, many studies explore the use of foundation models to solve challenges like long-horizon problems and effectively exploring and designing reward functions.
Norman (di2023towards) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. ROBOFUME (yang2024robot) employs a fine-tuned VLM as the sparse reward function for the RL algorithm, tackling the issue of the extensive human supervision needed for training or fine-tuning a policy in the real world. Eureka (ma2023eureka) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (xie2023text2reward) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (ye2023foundation) proposes using knowledge from foundation models as policy prior knowledge to improve sampling action efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
@@ 8 Manipulation Data Generation
Demonstration data plays a crucial role in robotic manipulation, particularly in the context of imitation learning (RTX). A common approach for gathering such demonstrations is human teleoperation in the real world. However, collecting real-world data often requires significant human labor and specialized teleoperation equipment. Recently, there has been a growing number of excellent developments in low-cost teleoperation hardware, which enables the collection of high-quality demonstration data (fang2024airexo; cheng2024open).
@@
To collect data in real environments, human effort is required for scene setup and data annotation (sermanet2023robovqa). There are currently two methods for data collection: the bottom-up approach and the top-down step-by-step approach. The bottom-up approach focuses on selecting a task to perform based on the current scene. Then, it uses methods like crowd-sourcing to label the data. The top-down approach involves a step-by-step process where decision-makers assign task labels and manage overheads, such as resets and scene preparations (sermanet2023robovqa). The robot then performs tasks according to these labels. RoboVQA (sermanet2023robovqa) shows that the bottom-up approach is more efficient in data collection compared to the top-down step-by-step approach. DIAL (xiao2022robotic) uses a fine-tuned CLIP to replace humans in labeling robot trajectories during bottom-up data collection. This transforms the robot manipulation dataset on the internet into the robot-language manipulation dataset. PAFF (ge2023policy) points out that incorrect robot trajectories can be linked to new tasks and uses fine-tuned CLIP to label the incorrect robot trajectories with appropriate task labels. The above methods demonstrate that high-level cognitive models can assist in data annotation. SOAR (zhou2024autonomous) shows that integrating a high-level cognitive model with a low-level control policy can result in a fully autonomous data-collection system in varied real-world environments.
@@
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (wang2023gensim). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. matas2018sim trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge.
@@
Compared to single-frame images and language data on the internet, internet videos contain information on the physics and dynamics of the world, as well as on human behaviors and actions (chandrasegaran2024hourvideo). This information is precisely what is required for manipulation tasks. Therefore, in this section, we also introduce the internet-scale video data for robot learning.
@@
Aloha(zhao2023learning) Online Teleoperation Joint Copy Third-person View \$2000(include robot) Dual-Arm; 2F Gripper
@@
GELLO(wu2023gello) Online Teleoperation Joint copy Third-person View \$300 Dual-Arm; 2F Gripper
@@
Human Plus(fu2024humanplus) Online Teleoperation End-effectors mapping Third-person View \$30(A RGB Camera) Humanoid Robot; Dexterous Hand
@@
Transteleop(li2020mobile) Online Teleoperation End-effectors mapping Third-person View \$170 Dual-Arm; Dexteours Hand
@@
OPEN TEACH(iyer2024open) Online Teleoperation End-effectors mapping First-person View \$500 Dual-Arm; Dexteours Hand/2F Gripper;
@@
Open-TeleVision(cheng2024open) Online Teleoperation End-effectors mapping First-person View \$3499 Humanoid Robot; Dexteours Hand
@@
AirExo (fang2024airexo) Offline Teleoperation Joint copy - \$600 Dual-Arm; 2F Gripper
@@
UMI(chi2024universal) Offline Teleoperation End-effectors mapping - \$371 Dual-Arm; 2F Gripper
@@
DexCap(wang2024dexcap) Offline Teleoperation End-effectors mapping - \$4000 Dual-Arm; Dexterous Hand
@@
VideoDex(shaw2023videodex) Offline Teleoperation End-effectors mapping - - Single-Arm; Dexterous Hand
@@ 8.1 Low-cost Teleoperation Device
The current low-cost teleoperation can be categorized into two types: online teleoperation and offline teleoperation. The distinction is similar to the difference between SLAM and SFM. Online teleoperation is a closed-loop interaction between a demonstrator and a robot (darvish2023teleoperation). In the forward process, human motion is measured using devices that combine various sensors, such as vision, IMUs, or multi-joint encoders. The motion data from the demonstrator is then retargeted to the robot’s space. This allows the robot to accurately follow the demonstrator’s demonstrated trajectory. During the backward process, sensor data from the robot, such as forces, torques, and tactile information, should be retargeted to the demonstrator’s space. As a result, the demonstrator can experience an immersive teleoperation environment by sensor data feedback. At the same time, the synchronization and real-time performance between the forward and backward processes are also crucial (darvish2023teleoperation). Offline teleoperation remove the reliance on real robots during data collection compared to online teleoperation (chi2024universal). Demonstrators directly perform tasks using handheld or wearable devices (fang2024airexo; chi2024universal; wang2024dexcap) or using cameras to record the task execution process (shaw2023videodex). They do not need to supervise real robots to complete the tasks and operate tasks using human’s direct view perspective. Therefore, offline teleoperation lacks the backward feedback process. Without relying on real robots, the devices become more portable and intuitive. However, this increases the precision requirements for the retargeting algorithm.
@@
The differences among current low-cost teleoperation devices lie primarily in two aspects. One is human motion measurement on both online teleoperation and offline teleoperation. The other is visual feedback on online teleoperation. Human motion measurement component can be broadly categorized into two classes: one aimed at capturing and mapping the pose of end-effectors (cheng2024open; liu2022robot; fu2024humanplus; chi2024universal; li2020mobile), and one exploited devices for joint copy (zhao2023learning; wu2023gello; fang2024airexo). Visual feedback can be generally classified into two types third-person view and first-person view (cheng2024open). The third-person view shows the demonstrator from an external position, offering a broader perspective of surroundings. In contrast, the first-person view mimics the robot’s perspective, providing an immersive and realistic experience such as teleoperation with VR/AR headset.
@@
For approaches capturing and mapping the pose of end-effectors, the common low-cost capturing devices include SpaceMouse (liu2022robot; zhu2023viola), cameras (cheng2024open; fu2024humanplus; iyer2024open; shaw2023videodex; li2019vision; fang2020vision), VR controllers (de2021leveraging; nakanishi2020towards) and IMU sensors (chi2024universal; li2020mobile; fang2017novel; fang2017robotic). The SpaceMouse based method passes the position and orientation of the SpaceMouse as action commands of end-effectors. This method is low-cost, easy operation, and easy implementation, but it is not suitable for dual-arm operations. In contrast, methods based on cameras and VR are well suited for bimanual teleoperation and VR offers the advantage of visual feedback compared to cameras. However, teleoperation methods based on cameras and VR heavily relies on the accuracy of pose estimation algorithms and often affected by occlusion (fu2024humanplus; pavlakos2024reconstructing; iyer2024open; cheng2024open). The main advantage of teleoperation devices based on IMU sensors lies in their wearability (li2020mobile; chi2024universal; wang2024dexcap). Due to this advantage, UMI (chi2024universal) and DexCap (wang2024dexcap) develope wearable devices capable of in-the-wild teleoperation and offline data collection.
@@
Above systems work in cartesian space, which needs inverse kinematic (IK) solver and off-the-shelf IK often suffering from fails when operating near singularities of the robot. Although some bilateral teleoperation systems use haptic feedback to provide a tangible sense of the robot’s kinematic constraints, they do not address the challenges of very tight operational spaces (silva2009phantom). Therefore, multi-joint encoder teleoperation devices can solve the IK problem by working in the joint space. The current design of multi-joint teleoperation devices is mainly divided into isomorphic and non-isomorphic devices (wu2023gello). Isomorphic devices refer to teleoperation systems using standard servo-based robotic arms to control manipulators with similar size and kinematics (zhao2023learning), while non-isomorphic devices use such arms to control manipulators with different size and kinematic properties. Non-isomorphic devices use kinematically equivalent structures based on DH parameters to map joint spaces between different properties (wu2023gello). AirExo (fang2024airexo) expands this low-cost and scalable platform into a wearable device to collect cheap in-the-wild demonstrations at scale.
@@
As for teleoperation visual feedback, most of methods (liu2022robot; zhu2023viola; fu2024humanplus; li2020mobile; zhao2023learning; wu2023gello) are use third side view that observe the robot task with the operator’s own eyes directly. However, this observation involves some visual errors. For example, there may be inaccuracies in the distance between the gripper and the object being manipulated. While for first-person view, due to wearing VR head (cheng2024open; iyer2024open; de2021leveraging; nakanishi2020towards), it allows operators to perceive the robot’s surroundings immersively. However, long time to use VR headset can lead to fatigue.
@@ 8.2 Simulator
The current mainstream simulators (zhou2023language) include PyBullet (coumans2016pybullet), MuJoCo (todorov2012mujoco), CoppeliaSim (rohmer2013v), NVIDIA Omniverse and Unity. Pybullet is easy to use and integrate, but its graphics are quite basic. It is not suitable for applications that require complex visual effects. Therefore, Pybullet is often used together with Blender (shi2024asgrasp). Mujoco offers a high-precision physics engine. It is suitable for simulating articulated and deformable object manipulation. However, it has a high entry barrier for beginners. CoppeliaSim offers a wide range of ready-made environments, objects, and prototyping robotic systems for users. However, when dealing with many robots or complex scenes, CoppeliaSim may encounter performance issues. NVIDIA Omniverse provides real-time physics simulation and realistic rendering. However, it requires significant computational resources. NVIDIA Omniverse offers many interfaces. Users can use these to develop various applications. For example, Issac Gym is a platform for robot reinforcement learning, developed using Omniverse. Unity offers rich visual effects and a user-friendly interface. It allows for the creation of highly interactive applications. However, its physics engine is still not precise enough. The basic components of a simulator are the physics engine and the renderer. Improvements in these components can enhance the capability of sensors in simulations, such as optical tactile sensors (chen2023tacchi). Learning-based simulators also show great potential. For example, Sora (videoworldsimulators2024) and UniSim (yang2023learning) use vast amounts of data from the internet to simulate the visual effects of many different actions.